SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (3rd July, 2007)

Attendees: Debbie Bard, Rainer Bartoldus, Kim Hojeong, Rahmat Rahmat, Mark Tibbetts.

Operations

  • Power glitch Friday 6am in EH turned off some TRG crates. Was paged, instructed pilot how to turn them back on.
  • Friday afternoon IR2 lost communication with PEP - in particular, the timing fiducial was lost. Had to powercycle TRG crates to reset them.
  • Hojeong did not recieve the email from Friday's page, despite being on the contact experts list.

EMT

  • Friday problems showed up as EMT cable errors in the fastmon plots.
  • Mark is now on the babatrg forwarding list.

L3

  • Arik will write L3 tracking BAD note but someone is still needed to finish the MC studies.

Data Quality and Monitoring

  • Run 6 TRG stripcharts can be found on the QA webpage: here.
  • Rahmat still has no account with Padova. He will contact them again and post to Padova SOS hypernews for ideas about the possibly different setup there.

Trigger Configuration

  • Hojeong had troubles running L1TNtupleKanApp and will post to hypernews if they persist.

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.